Grade 6General Physics

Asolid of cylinder has radius of 2cm and a length of 7cm .it has adensity of 3.1g/cm what is the mass of the cylinder ?

Hadia
7 Years agoGrade 6

1 Answer

Arun
7 Years ago
Dear student
 
mass = volume * density
 
 = pi * r^2 * h * density
 
 = 22/7 * 4 * 7 * 3.1 = 272.8 gm
 
Regards
Arun